OpenStudy (anonymous):

Solve x2 - 8x - 9 = 0. Rewrite the equation so that it is of the form x2 + bx = c.

OpenStudy (welshfella):

just add 9 to both sides of theequation

OpenStudy (anonymous):

I made a mistake in the question. x^2 - 8x - 9 = 0 and x^2 + bx = c.

OpenStudy (welshfella):

what do you get when you add 9 ?

OpenStudy (welshfella):

x^2 - 8x - 9 = 0 x^2 - 8x - 9 + 9 = 0 + 9